A cylindrical tank has a height of 6.0 ft and a diameter of 3.0 ft. What is its capacity in liters if the volume of a cylinder can be found by the equation 2π r 2 h? Use: 3.28 ft = 1.0 m; 1000 L = 1 m3
1
Expert's answer
2020-10-05T10:52:14-0400
"V=\pi r^2h\\V=\pi\left(\frac{1.5}{3.28}\right)^2\left(\frac{6}{3.28}\right)=1.202\ m^3=1202\ L"